Assume that a certain location on the Earth reflects 29.0% of the incident sunlight from its clouds and surface.

(a) If the intensity of solar radiation is 1373 W/m2, what is the radiation pressure (Prad) on the Earth at this location? ________ Pa

(b) Compare this to normal atmospheric pressure (Pa) at the Earth's surface, which is 101 kPa.

Prad is ________ smaller than atmospheric pressure
Answer #1

a) Prad= (Ef*cos2\alpha)/c (N/M2 or Pa)

= (1353/(3*108)) *cos229

=0.2523*10-5 Pa

b) 1Pa=101kPa

then 0.2523*10-5Pa= 0.25*10-3 kPa

so, Prad is 10-2 is smaller the atmospheric pressure.